At an air show, a jet plane has velocity componentsvx = 695km/h and vy = 485km/h at time 4.05sand vx = 848km/h and vy = 345km/h at time 6.52s .

a.For this time interval, find the x component of the plane's average acceleration.

b.For this time interval, find the

y component of the plane's average acceleration.

c.For this time interval, find the magnitude of its average acceleration.

d.For this time interval, find the direction of its average acceleration.
Answer #1


A) ax = (848-695)*(5/18)/(6.52-4.05) = 17.2 m/s^2
B)ay = (345-485)*(5/18)/(6.52-4.05) = -15.75 m/s^2
C) a = sqrt(ax^2+ay^2) = sqrt(17.2^2+15.75^2) = 23.32 m/s^2

D) theta = arctan(ay/ax) = arc tan( -15.75/17.2) = -42.48 degrees from the +X-axis